# Estimation of partially known Gaussian graphical models with score-based structural priors

### 25 Jan 2024 v1

@@ F.2 Proof of Lemma 2
G EXPERIMENTAL DETAILS
G.1 Hyperparameters
In terms of the Langevin sampler, in all the experiments we use $L=10$ noise levels, evenly spaced between $\sigma_{1}=0.5$ and $\sigma_{L}=0.03$, and $T=300$ steps per level. We set the step size at $\epsilon=10^{-6}$.
In order to tune the regularization parameter $\lambda$ for graphical lasso, we fit a function of the form $\lambda(k)=a\log(k)^{2}+b\log(k)+c$ using a training set, and then evaluate that function at inference time for the given value of $k$. We selected this specific functional form as it showed a very satisfactory fit for all of our cases. An example is shown in Figure SM3.
Figure SM3: Fitting of $\lambda(k)=a\log(k)^{2}+b\log(k)+c$ for the ego-nets dataset with $|{\mathcal{U}}|=0.5\dim({{\mathbf{a}}})$. The orange curve on the right is the one used at inference time.
The thresholds used to obtain a binary-valued estimate of ${\mathbf{A}}_{0}$ when using GraphSAGE, the thresholding of $\hat{\boldsymbol{\Theta}}$ and our own method (in this last case, the threshold is $\tau_{k}$) are set using the same procedure. Given $R$ simulations (in all of our experiments, $R=100$), we take $R/2$ of those continuous estimated matrices and find the threshold within some grid ${\mathcal{T}}$ that works the best for each method. Then, with those tuned values, we threshold the other $R/2$ matrices left and only evaluate performance over them. All the plots shown in Section 4 from the original paper as well as in this document correspond to averages following this procedure $S$ times (i.e., across $S=10$ different train/test splits in our case), each with its own threshold. A more clear description of our method is provided in Algorithm SM2.
Algorithm SM2 Metric computation
@@ G.1 Hyperparameters
G.2 Datasets’ details
For the grids, we generated graphs with $5040\!\leq\!n\!\leq\!$ 50. Additionally, to introduce randomness into the graphs, we added between $2$ and $5$ edges uniformly at random to the edge set ${\mathcal{E}}$. We trained a GNN with $|{\mathcal{A}}|=5000$ graphs generated with the described procedure.
@@ G.2 Datasets’ details
G.3 Implementation details
All experiments shown both in the main paper as well as in this document, together with the training of the EDP-GNNs, were run on an NVIDIA DGX A100 system. Our code is provided as part of the supplementary material and not in a GitHub repository for the sake of anonymity. The code corresponding to the EDP-GNN implementation was taken from the original repository of Niu et al., (2020) under a GPL-3.0 license.
Regarding the computational complexity, the key steps to consider from Algorithm 1 are those in lines 3, 12, and 13. The computation time of $\hat{\boldsymbol{\Theta}}$ is ${\mathcal{O}}(n^{3})$ for dense problems, but much less in sparse ones (Friedman et al.,, 2008), and even less if some of the entries are fixed as is the case of (3.1). However, here we consider the worst-case scenario. After computing $\hat{\boldsymbol{\Theta}}$, we evaluate (27), which requires to compute the inverse of $\tilde{\boldsymbol{\Theta}}$, leading to a time complexity of ${\mathcal{O}}(n^{3})$ as well. Feed-forwarding the GNN consists of just matrix and vector multiplications that scale as ${\mathcal{O}}(n^{2})$. Therefore, the dominating steps are the first two. The step in 12 is repeated $L\cdot T$ times per sample, and since the sampling of each ${\mathbf{A}}_{i}$ is completely parallelizable, the total complexity of Algorithm 1 is ${\mathcal{O}}(L\cdot T\cdot n^{3})$, considering that typically $LT\gg 1$.
H ADDITIONAL EXPERIMENTS
In this section, we provide more experiments to gain additional insights on our method.
H.1 Performance dependence on $|{\mathcal{A}}|$
@@ H ADDITIONAL EXPERIMENTS
$p({\mathbf{A}})=\frac{1}{C_{\psi}(\boldsymbol{\beta})}\exp\left(\boldsymbol{\beta}^{\top}\boldsymbol{\psi}({\mathbf{A}})\right).$ (SM39)
@@
$\boldsymbol{\psi}({\mathbf{A}})=\begin{bmatrix}\mathrm{AKS}_{\gamma}({\mathbf{A}})&\frac{1}{2}\sum_{ij}A_{ij}\end{bmatrix}^{\top}.$ (SM40)
@@
$\mathrm{AKS}_{\gamma}({\mathbf{A}})=\sum_{d=2}^{p-1}(-1)^{d}\frac{S_{d}({\mathbf{A}})}{\gamma^{d-2}},$ (SM41)
with $S_{d}(\cdot)$ being the number of $d$-stars in the given graph and $\gamma$ a constant. In our simulations we set $\gamma=0.3$. The second statistic corresponds to the number of edges $|{\mathcal{E}}|$. We use $\boldsymbol{\beta}=\begin{bmatrix}0.7&-2\end{bmatrix}^{\top}$ as coefficients. The graphs we generate have $n=50$ nodes, and we remove $|{\mathcal{U}}|=30$ values from ${\mathbf{A}}_{0}$. We drop zeros and ones with equal probability, so that we can use accuracy instead of F1 score for this experiment. We generated $|{\mathcal{A}}|=1000$ graphs from this distribution to train the EDP-GNN and, to evaluate the impact the dataset size $|{\mathcal{A}}|$ has on the estimation performance, we compare the edge prediction accuracy when fewer training samples are used. The reported accuracies correspond to the average over $25$ runs, following the same procedure described in Section G.1.
Figure SM4: Prediction accuracy of several methods using ERGM graphs with $|{\mathcal{U}}|=30$. The comparison includes five versions of Algorithm 1, each considering a prior that was learned with datasets ${\mathcal{A}}$ of different sizes (the value of $|{\mathcal{A}}|$ is indicated in the legend), as well as three benchmarks available in the literature.
Figure SM4 shows that the accuracy improves drastically as $|{\mathcal{A}}|$ increases, as expected. However, this enhancement seems to saturate as the dataset size approaches $|{\mathcal{A}}|=1000$. All the other GGM estimation methods perform similarly, logically presenting a higher prediction accuracy as $k$ increases. Notably, whenever $|{\mathcal{A}}|\geq 300$, all other methods underperform our algorithm .
GraphSAGE performs rather poorly – it predicts 00 or $1$ uniformly at random. Additionally, the incorporation of additional observations does not increase its accuracy. As discussed in detail in the main paper, this behavior is expected and associated with the fact of GraphSAGE aggregating data using local neighborhoods, which is not a good fit for the setup at hand.
H.2 Performance dependence on $|{\mathcal{U}}|$
@@
In this experiment, we use the AUC score in lieu of F1 to assess the prediction performance. This allows us to avoid the tuning of $\tau_{k}$, which is not needed for this analysis since we are comparing three versions of Algorithm 1 that return continuous predictions if no thresholding is applied. Results are shown in Figure SM5.
The main observations are: i) the information provided by ${\mathcal{A}}$ decreases as $|{\mathcal{U}}|$ increases; and ii) LPost always outperforms LPr and LL. Both are expected results that were observed in previous experiments. Analyzing more specific details, we note that the performance of LPr drops rapidly as $|{\mathcal{U}}|$ increases, while that of LPost decays more slowly. The same is true for LL, whose AUC levels are a quite stable and decay at a very slow pace. Finally, it is worth noticing that when a large portion of the graph is unknown, LPost approaches LL, since the information provided by $|{\mathcal{A}}|$ becomes less useful.
Figure SM5: AUC score achieved by different versions Algorithm 1 when using an ego-net prior and $k=100$ observations. The horizontal axis represents different values of $|{\mathcal{U}}|$. For this experiment we skip the thresholding implemented in line 25 of Algorithm 1.

### 23 Feb 2024 v3

@@ A.2 Proof of Lemma 2
B EXPERIMENTAL DETAILS
B.1 Hyperparameters
Regarding the Langevin sampler, in all the experiments, we use $L=10$ noise levels, evenly spaced between $\sigma_{1}=0.5$ and $\sigma_{L}=0.03$, and $T=300$ steps per level. We set the step size at $\epsilon=10^{-6}$.
To tune the regularization parameter $\lambda$ for the graphical lasso, we fit a function of the form $\lambda(k)=a\log(k)^{2}+b\log(k)+c$ using a training set and then evaluate that function at inference time for the given value of $k$. We selected this specific functional form as it showed a very satisfactory fit for all of our cases. An example is shown in Figure 5.
Figure 5: Fitting of $\lambda(k)=a\log(k)^{2}+b\log(k)+c$ for the ego-nets dataset with $|{\mathcal{U}}|=0.5\dim({{\mathbf{a}}})$. The orange curve on the right is the one used at inference time.
The thresholds used to obtain a binary-valued estimate of ${\mathbf{A}}_{0}$ when using GraphSAGE, the thresholding of $\hat{\boldsymbol{\Theta}}$ and our method (in this last case, the threshold is $\tau_{k}$) are set using the same procedure. Given $R$ simulations (in all of our experiments, $R=100$), we take $R/2$ of those continuous-valued matrices and find the threshold within some grid ${\mathcal{T}}$ that works the best for each method. Then, with those tuned values, we threshold the other $R/2$ matrices left and only evaluate performance over them. All the plots shown in Section 4 from the original paper, as well as in this document, correspond to averages following this procedure $S$ times (i.e., across $S=10$ different train/test splits in our case), each with its threshold. Algorithm 2 provides a more precise description of our method.
Algorithm 2 Metric computation
@@ B.1 Hyperparameters
B.2 Datasets’ details
For the grids, we generated graphs with $5040\!\leq\!n\!\leq\!$ 50. Additionally, to introduce randomness into the graphs, we uniformly added between $2$ and $5$ edges at random to the edge set ${\mathcal{E}}$. We trained a GNN with $|{\mathcal{A}}|=5000$ graphs generated with the described procedure.
@@ B.2 Datasets’ details
B.3 Implementation details
All experiments shown in the main paper and in this document and the training of the EDP-GNNs were run on an NVIDIA DGX A100 system. Our code is provided as part of the supplementary material and not in a GitHub repository for the sake of anonymity. The code corresponding to the EDP-GNN implementation was taken from the original repository of Niu et al., (2020) under a GPL-3.0 license.
Regarding the computational complexity, the key steps to consider from Algorithm 1 are those in lines 3, 12, and 13. The computation time of $\hat{\boldsymbol{\Theta}}$ is ${\mathcal{O}}(n^{3})$ for dense problems, but much less in sparse ones (Friedman et al.,, 2008), and even less if some of the entries are fixed as is the case of (3.1). However, here, we consider the worst-case scenario. After computing $\hat{\boldsymbol{\Theta}}$, we evaluate (27), which requires to compute the inverse of $\tilde{\boldsymbol{\Theta}}$, leading to a time complexity of ${\mathcal{O}}(n^{3})$ as well. Feed-forwarding the GNN consists of just matrix and vector multiplications that scale as ${\mathcal{O}}(n^{2})$. Therefore, the dominating steps are the first two. The step in 12 is repeated $L\cdot T$ times per sample, and since the sampling of each ${\mathbf{A}}_{i}$ is completely parallelizable, the total complexity of Algorithm 1 is ${\mathcal{O}}(L\cdot T\cdot n^{3})$, considering that typically $LT\gg 1$.
C ADDITIONAL EXPERIMENTS
This section provides more experiments to gain additional insights into our method.
C.1 Performance dependence on $|{\mathcal{A}}|$
@@ C ADDITIONAL EXPERIMENTS
$p({\mathbf{A}})=\frac{1}{C_{\psi}(\boldsymbol{\beta})}\exp\left(\boldsymbol{\beta}^{\top}\boldsymbol{\psi}({\mathbf{A}})\right).$ (41)
@@
$\boldsymbol{\psi}({\mathbf{A}})=\begin{bmatrix}\mathrm{AKS}_{\gamma}({\mathbf{A}})&\frac{1}{2}\sum_{ij}A_{ij}\end{bmatrix}^{\top}.$ (42)
@@
$\mathrm{AKS}_{\gamma}({\mathbf{A}})=\sum_{d=2}^{p-1}(-1)^{d}\frac{S_{d}({\mathbf{A}})}{\gamma^{d-2}},$ (43)
with $S_{d}(\cdot)$ being the number of $d$-stars in the given graph and $\gamma$ a constant. In our simulations, we set $\gamma=0.3$. The second statistic corresponds to the number of edges $|{\mathcal{E}}|$. We use $\boldsymbol{\beta}=\begin{bmatrix}0.7&-2\end{bmatrix}^{\top}$ as coefficients. The graphs we generate have $n=50$ nodes, and we remove $|{\mathcal{U}}|=30$ values from ${\mathbf{A}}_{0}$. We drop zeros and ones with equal probability so that we can use accuracy instead of F1 score for this experiment. We generated $|{\mathcal{A}}|=1000$ graphs from this distribution to train the EDP-GNN and, to evaluate the impact the dataset size $|{\mathcal{A}}|$ has on the estimation performance, we compare the edge prediction accuracy when fewer training samples are used. The reported accuracies correspond to the average over $25$ runs, following the procedure described in Section B.1.
Figure 6: Prediction accuracy of several methods using ERGM graphs with $|{\mathcal{U}}|=30$. The comparison includes five versions of Algorithm 1, each considering a prior that was learned with datasets ${\mathcal{A}}$ of different sizes (the value of $|{\mathcal{A}}|$ is indicated in the legend), as well as three benchmarks available in the literature.
Figure 6 shows that the accuracy improves drastically as $|{\mathcal{A}}|$ increases, as expected. However, this enhancement seems to saturate as the dataset size approaches $|{\mathcal{A}}|=1000$. All the other GGM estimation methods perform similarly, logically presenting a higher prediction accuracy as $k$ increases. Notably, whenever $|{\mathcal{A}}|\geq 300$, all other methods underperform our algorithm.
GraphSAGE performs poorly – it predicts 00 or $1$ uniformly at random. Additionally, the incorporation of additional observations does not increase its accuracy. As discussed in detail in the main paper, this behavior is expected and associated with GraphSAGE aggregating data using local neighborhoods, which is not a good fit for the setup at hand.
C.2 Performance dependence on $|{\mathcal{U}}|$
@@
We use the AUC score instead of F1 in this experiment to assess the prediction performance. This allows us to avoid the tuning of $\tau_{k}$, which is unnecessary for this analysis since we are comparing three versions of Algorithm 1 that return continuous predictions if no thresholding is applied. Results are shown in Figure 7.
The main observations are: i) the information provided by ${\mathcal{A}}$ decreases as $|{\mathcal{U}}|$ increases, and ii) LPost consistently outperforms LPr and LL. Both are expected results that were observed in previous experiments. Analyzing more specific details, we note that the performance of LPr drops rapidly as $|{\mathcal{U}}|$ increases, while that of LPost decays more slowly. The same is true for LL, whose AUC levels are stable and decay slowly. Finally, it is worth noticing that when a large portion of the graph is unknown, LPost approaches LL since the information provided by $|{\mathcal{A}}|$ becomes less valuable.
Figure 7: AUC score achieved by different versions Algorithm 1 when using an ego-net prior and $k=100$ observations. The horizontal axis represents different values of $|{\mathcal{U}}|$. For this experiment we skip the thresholding implemented in line 25 of Algorithm 1.